Daniel Arvesen

12.1.03

Greek and Roman Mythology

Paper 3

Heroes and the Underworld

Vladimir Propp, the famous Russian Formalist, noticed in Russian folklore that each tale followed a linear structure. Classical scholars have then applied Propp’s theory to Greek mythology and discovered that many of the structural points of his theory apply to the myths of Greek heroes. Nine goals have been described that Greek heroes have in common when they go on their heroic quest. (Morford 14) The second to last stop is a conquest against death in which the hero travels to the Underworld and then returns. Only a handful of characters in Greek mythology have accomplished this step and those that have stand out amongst the Greek heroes.

Although the descent to the Underworld is an easy task, retracing your steps and getting out is nearly impossible and only a certain few have succeeded in doing so; those that have are the ones that many consider to be heroes. Another reason that this is significant is that descending to the Underworld and returning is a conquest of the afterlife. The Greeks feared the afterlife because it is a stark difference from the norma life that the Greeks were used to. The afterlife and Underworld was a dusty, fiery, muddy world with nothing to hold on to. (Hillman 40)

Propp’s theory can be applied to many of the myths of classical mythology, including the myths of Heracles, Theseus, Jason, and Perseus. Of these, only Heracles and Theseus visited the underworld, but all four share the common elements that Propp noticed. Not included in this group are Odysseus and Orpheus, two mythical figures who went on a quest that did not follow Propp’s theory but visited the Underworld and were later recognized as heroes.

Conquering death is basically doing the unthinkable. Greek men and women usually never considered visiting the Underworld. They were terrified of it and did not want to think about it. Those that were brave enough to enter the Underworld are those that many people consider to be heroes.

Arguably the most famous hero in Greek mythology would be Heracles. Hera had made Heracles go mad in which he killed Megara and her children. He was told by a priestess of Apollo that he must go to Tiryns for twelve years and serve Eurystheus and perform all the deeds that he imposed on him. If he could perform all of them then he would become immortal. Of the twelve labors that he performed, only the last three are conquests of death. These were The Cattle of Geryon, The Apples of the Hesperides, and Cerberus.

No mortal man had ever dared to fight Geryon and it is for this reason that Eurystheus chose Heracles to perform this task. Heracles sailed to the western land of Erythia where Geryon is located, killed him, and then returned to Eurystheus. In this deed Heracles does not actually visit the Underworld, but he does conquer death because he killed the undefeated Geryon.

His next quest was to get the Apples of the Hesperides. The Hesperides were three daughters of Night who lived far in the west and guarded a tree that grew golden apples. He was told by Nereus to go to the far land where Atlas lives. There he found the apples of the Hesperides, but he was told by Prometheus, who he freed on the way there, that Atlas should pick them instead. So Heracles takes the weight of the heavens on his shoulders while Atlas picks the apples. Heracles then tricks Atlas into bearing the weight of the heavens again, and Heracles takes the apples and leaves.

There are three reasons why this deed is a conquest of death. The obvious reason is that the apples are symbols of immortality since they were a gift given by Gaea at Zeus and Hera’s wedding. The second reason is that, according to Morford and Lenardon, the tree that the apples grew on is a sort of Tree of Life. (Morford 528) The third and less obvious reason is that all of the deeds that Eurystheus imposed on Heracles were meant to kill him and so far he has survived.

Heracles’ final deed was to fetch Cerberus, the three-headed hound of Hades. This deed is very obviously a conquest of death. Heracles entered the Underworld in the city of Taenarus in Laconia and was accompanied by Hermes. (Schwab 180) In the Underworld Heracles sees Theseus and Pirithous chained for their attempt to abduct Persephone. Heracles releases Theseus but does not bother freeing Pirithous. Upon coming to the gates of Hades Heracles fires an arrow at Hades that pierces his shoulder. With Hades in pain Heracles asks to take Cerberus. Hades allows him but imposes the condition that he must master the dog without using any of his weapons. With this condition, his difficult quest got even harder. Heracles then subdued Cerberus by wrestling him, left through another entrance near Troezen in Argolis and brought Cerberus to Eurystheus. Upon seeing Cerberus, Eurystheus reverses Heracles’ fate, and then Heracles returns the dog to Hades.

While reading about Heracles, one notices that although he is very strong and powerful, he still receives help during his quest. In the story of the Apples of the Hesperides Heracles receives help from Nereus, Atlas, and Prometheus. This shows that although he is very strong, there is still a human element within him. He is not all-powerful therefore he has to look for help when he needs it.

The story of Theseus differs from that of Heracles because he chooses to go to go on his heroic quest whereas Heracles was forced by Eurystheus to perform the twelve deeds. Theseus does not actually visit the Underworld until he is considered a hero. Theseus and his friend Pirithous both wanted a wife so Theseus decided to abduct Helen of Troy and Pirithous wanted Persephone, the wife of Hades. Theseus and Pirithous then descended to the Underworld but were caught by Hades and locked up in magic chairs. It was not until Heracles came while trying to get Cerberus that Theseus was freed. Pirithous, however, remained locked in the chair.

Theseus’ story is unusual because it is actually Heracles that conquers death and not Theseus. He is the one that frees Theseus, and therefore Theseus is associated with Heracles and his conquest of death. However, because Theseus entered the Underworld and returned, his heroic status is enhanced even more.

Theseus and Pirithous’s reasons to enter the Underworld are very different than those of the other heroes. Clearly they knew the dangers of entering the Underworld and the risk that they are taking. However, love conquers all in this case, and they will go to all ends to get the woman that they desire.

Orpheus is unlike other Greek heroes in that he did not achieve his status by fighting monsters or in war. He is more of an “everyman’s” hero. He is a poet and a singer, not a warrior or king. Because of this, his descent to the Underworld was unusual and the circumstances that brought him there make him stand out against other heroes.

The story of Orpheus is another of love conquers all. Orpheus had just married Eurydice when she is bitten on her ankle by a serpent and dies. Orpheus decides to descend into the Underworld and retrieve her. Likes Heracles, he enters the Underworld through the entrance in Taenarus. He approaches Hades and sings to him while plucking his lyre:

O deities of the world below the earth, into which all of us who are mortal return, if it is right and you allow me to utter the truth, laying aside evasion and falsehood, I did not come down to see the realms of Tartarus or to bind the triple neck, bristling with serpents, of the monstrous hound descended from Medusa; the cause of my journey is my wife; she stepped on a snake, and its venom coursing through her veins stole from her the bloom of her years. I wanted to be able to endure, and I admit that I have tried; but Love has conquered. He is a god who is well known in the world above; I suspect that he is famous even here as well (although I do not know for sure); if the story of the abduction of long ago is not a lie, Love also brought you two together. (Metamorphosis 10. 1-64)

Upon completion of his song, everyone in the Underworld weeps, and Hades and Persephone, being touched by Orpheus’ song, call for Eurydice. (Metamorphosis 10. 50-52) Eurydice is allowed to return as long as Orpheus does not turn back his gaze until he has left the Underworld. They manage to get very close to the top, but Orpheus, worried that she might not be well, “through love he turned and looked, and with his gaze she slipped away and down.” (Metamorphosis) Thus, Eurydice dies again and Orpheus is shocked.

Orpheus’s song is very striking in that he raises many points about the Underworld and love. He points out that Love extends even to the depths of the Underworld and that “Love also brought you two together” [Hades and Persephone] (Metamorphosis). He explains that Love has conquered him and that is his reason for descending to the Underworld.

Odysseus is very similar to Jason and Perseus in that they are all considered heroes but none of them visit the Underworld. Odysseus stands out, however, in that he talks to the ghosts of Hades but he does not actually visit the Underworld.

Odysseus is told by Circe that to find out how he can get home he must question Tiresias. To do that he must go to the land of the Cimmerians, which is in the ”outermost Ocean“ (Odyssey 11. 14). There he performed a ritual of slitting sheep’s throats over a pit filled with seasonings. Soon the souls of the dead arose and Odysseus fended them away from drinking the blood in the pit until he questioned Tiresias. Tiresias approaches and tells Odysseus that in order for the seas to be calm he must sacrifice a ram, a bull, and a boar to Poseidon.

By examining the myths of these four heroes, one gets an insight into Greek life. We see that Love extends even into the Underworld, that humans are not all-powerful, and that Love conquers all.

Bibliography

1. Graves, Robert. The Greek Myths. New York: Penguin Books, 1992.

2. Hillman, George. The Dream and the Underworld. New York: Harper & Row, Publishers, 1979.

3. Lombardo, Stanley, ed. The Essential Homer. Indianapolis: Hackett Publishing Company, Inc, 2000.

4. Melville, A.D. Metamorphosis. New York: Oxford University Press, 1986.

5. Morford, Mark P.O., Robert J. Lenardon. Classical Mythology. New York: Oxford University Press, Inc., 2003.

6. Schwab, Gustav. Gods & Heroes. New York: Pantheon Books, 1946.